CSC/CPE 580

Intelligent Agents

Spring 2002

Franz J. Kurfess

Computer Science Department California Polytechnic State University

San Luis Obispo

Course Overview

Introduction Agent Architectures Problem-Solving Agents Knowledge-Based Agents Planning Agents Learning Agents Communicating Agents Implementation Languages for Agents Information Filtering and Searching Applications • Business Applications

• Medical Applications

• Manufacturing Applications Agents as Reuse Methodology Mobility in Agents Multi-Agent Interaction Security